OpenStudy (anonymous):

find the length of the 30 swg nichrome wire of a diameter 0.315 mm that would be required to construct a heating coil of 1000 watt rating to be operated at 230 v supply. resistivity for nichrome can be assumed as 110 micro-ohm-cm

OpenStudy (anonymous):

@Pompeii00

OpenStudy (anonymous):

First , you have to find the resistanve, using the formula R= V^2/P, where V is voltage and P power. Having an answer of 52.9ohms. Secondly use the formula of resistance R= (Rho*4*L)/(Pid^2), whre Rho is resistivity, d is diameter, Pi is 3.14 and L is length. Making L subject of the formula, becomes L= (RPid^2)/(Rho*4) and replacing with given values you get an answer of 15m.